# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ Acknowledgment
J.Y.W. received funding from MEXT/JSPS KAKENHI Grant Numbers JP21K03357 and JP24H00001. H.O. acknowledges the support from the "Evolutionary Studies of Complex Adaptive Systems" Research Grant from the Research Center for Integrative Evolutionary Science, SOKENDAI. Y. K. received funding from MEXT/JSPS KAKENHI Grant Numbers 25K06712 and 25K00199. E.B. received funding from the Deutsche Forschungsgemeinschaft, Germany (DFG, German Research Foundation) — Project-ID 317210226 — SFB 1283 and thanks Tom Britton (Stockholm) for an enlightening discussion about the connection to epidemiology.
@@ A Sojourn times of the birth-death process
Consider our general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $[N]_{0}$ with unique absorbing state 0 and birth and death rates $\lambda_{j}$ and $\mu_{j}$ with $\lambda_{0}=0$, $\lambda_{j}\geq 0$ for $j\in[N-1]$, and $\mu_{j}>0$ for $j\in[N]$, complemented by $\mu_{0}=\lambda_{N}=0$.
@@
Let $Z(0)=i\in[N]$, let $T_{\ell}$ be the time where the process hits state $\ell\;(1\leq\ell\leq i)$ for the first time, and let $S_{j}^{(\ell)}$ be the total sojourn time in state $j\;(\ell\leq j\leq N)$ in the interval $[T_{\ell},T_{\ell-1}]$. We then have
$\mathbb{E}\big{[}S_{j}^{(\ell)}\mid Z(0)=i\big{]}=\frac{\lambda_{\ell}\cdots\lambda_{j-1}}{\mu_{\ell}\cdots\mu_{j}}=\mathrel{\mathop{:}}\eta_{\ell j},$
where the empty product is 1.
@@
The hitting times are finite almost surely due to the almost sure absorption in 0. Fix $j\in[N]$. We first consider $\ell=j$ and let $M_{j}$ be the number of times the chain visits $j$ before moving to $j-1$ (where the initial visit is also counted). When in $j$, the chain moves up and down with probability $\lambda_{j}/(\lambda_{j}+\mu_{j})$ and $\mu_{j}/(\lambda_{j}+\mu_{j})$, respectively, so $M_{j}$ equals $1$ plus a geometric random variable with parameter $\mu_{j}/(\lambda_{j}+\mu_{j})$, where a geometric random variable with success parameter $p$ refers to the number of trials before the first success; note that, for $j=N$, we have the limiting case of $p=1$ and hence $M_{N}=1$. So $\mathbb{E}\left[M_{j}\right]=1+\lambda_{j}/\mu_{j}=(\lambda_{j}+\mu_{j})/\mu_{j}$. Each such visit has a mean duration of $1/(\lambda_{j}+\mu_{j})$, so, by Wald’s identity,
$\mathbb{E}\left[(S_{j}^{(j)}\mid Z(0)=i\right]=\mathbb{E}\left[M_{j}\right]\frac{1}{\lambda_{j}+\mu_{j}}=\frac{1}{\mu_{j}}=\eta_{jj}.$ (A.1)
This proves the claim for $\ell=j$ and serves as the start for induction in $\ell$. For the step from $\ell$ to $\ell-1$, let $N_{\ell-1\to\ell}$ be the number of transitions from $\ell-1$ to $\ell$ in $[T_{\ell-1},T_{\ell-2}]$; then $N_{\ell-1\to\ell}$ follows the geometric distribution with parameter $\mu_{\ell-1}/(\mu_{\ell-1}+\lambda_{\ell-1})$, so $\mathbb{E}\left[N_{\ell-1\to\ell}\right]=\lambda_{\ell-1}/\mu_{\ell-1}$. Let further $S^{(\ell)}_{jn}$ be the sojourn time in $j$ in the interval
[time of the $[\text{time of the }n\text{th arrival in }\ell\text{ after }T_{\ell-1},\text{time of the following arrival in }\ell-1]$
for $n\in[N_{\ell-1\to\ell}]$; clearly, $S^{(\ell)}_{j1}$ equals $S^{(\ell)}_{j}$ in distribution, and, by the Markov property, the $S^{(\ell)}_{jn}$ are iid. Since $S^{(\ell-1)}_{j}=\sum^{N_{\ell-1\to\ell}}_{n=1}S^{(\ell)}_{jn}$, we thus get
$\mathbb{E}\left[S^{(\ell-1)}_{j}\right]=\mathbb{E}\left[N_{\ell-1\to\ell}\right]\mathbb{E}\left[S^{(\ell)}_{j}\right]=\frac{\lambda_{\ell-1}}{\mu_{\ell-1}}\mathbb{E}\left[S^{(j)}_{\ell}\right]=\frac{\lambda_{\ell-1}\cdot\ldots\cdot\lambda_{j-1}}{\mu_{\ell-1}\cdot\ldots\cdot\mu_{j}}=\eta_{\ell-1,j},$
where we have used Wald’s identity in the first step and the induction hypothesis in the penultimate one.
@@
Let us also remark that there is a nice intuitive argument for (A.1). If $\lambda_{j}=0$, the transition to $j-1$ at rate $\mu_{j}$ is the only way out of $j$, so the sojourn time in $j$ follows $\mathrm{Exp}(\mu_{j})$ and (A.1) is clear. For $\lambda_{j}>0$, the sojourn time in $j$ may be interrupted by one or several excursions to states $>j$, so the exponential clock for a death event out of state $j$ runs intermittently; but, due to the memoryless property of the exponential distribution, the total sojourn time in $j$ still follows $\mathrm{Exp}(\mu_{j})$, thus (A.1) remains true.
@@ B A derivation of the moments of the path integrals of birth-death processes
The moments (24) of the path integral (20) are usually derived via Laplace transforms (see, for example, Goel and Richter-Dyn (1974, App. D)). Here we take an alternative route that bypasses Laplace transforms and derives a first-step equation in a direct way, thus validating (62) derived in a heuristic way by Norden (1982) via the backward equation.
We first observe, following eq.(83) in Crawford et al. (2018), that our path integral (20) for the birth-death process $(Z(t))_{t\geqslant 0}$ restricted to $[N]_{0}$ with birth and death rates $\lambda_{j}\geq 0$ and $\mu_{j}>0$ for $j\in[N]$, $\lambda_{0}=0$, and the convention $\mu_{0}=\lambda_{N}=0$, equals the first-passge time
$T_{i}^{*}\mathrel{\mathop{:}}=\inf\{t:Z^{*}(t)=0\mid Z^{*}(0)=i\},\quad i\in[N]_{0},$ (B1)
of the “modified" birth-death process $(Z^{*}(t))_{t\geqslant 0}$ with birth and death rates
$\lambda_{j}^{*}\mathrel{\mathop{:}}=\frac{\lambda_{j}}{f(j)},\quad\mu_{j}^{*}\mathrel{\mathop{:}}=\frac{\mu_{j}}{f(j)}\quad\text{for }j\in[N],\quad\text{and}\quad\lambda_{0}^{*}=\mu_{0}^{*}=0$ (B2)
(this is well defined since $f(j)>0$ for $j\in[N]$). Note that, due to the finite state space, $T_{i}^{*}$ is almost surely finite.
@@
Let us now calculate the $m$-th moment of $T_{i}^{*}$, following the standard first-step approach outlined, for example, in (Pinsky and Karlin, 2011, Sec. 6.5.2). Suppose that we are currently in state $i\in[N]$. The sojourn time in this state, whose realisation is denoted by $r$ below, follows the exponential distribution with parameter $\lambda_{i}^{*}+\mu_{i}^{*}$. After the transition away from $i$, the process experiences a waiting time until absorption that is distributed as $T_{i+1}^{*}$ or $T_{i-1}^{*}$, depending on whether it moves from $i$ to $i+1$ (probability $\lambda_{i}^{*}/(\lambda_{i}^{*}+\mu_{i}^{*})$) or $i-1$ (probability $\mu_{i}^{*}/(\lambda_{i}^{*}+\mu_{i}^{*})$). Thus we obtain
@@
$\begin{split}\mathbb{E}\left[(T_{i}^{*})^{m}\right]&=\int_{0}^{\infty}\left\{\frac{\lambda_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i+1}^{*}+r)^{m}\right]+\frac{\mu_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i-1}^{*}+r)^{m}\right]\right\}(\lambda_{i}^{*}+\mu_{i}^{*})e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\\ &=\lambda_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i+1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]+\mu_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i-1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]\end{split}$ (B3)
@@
(products containing the factor $\lambda_{N}^{*}=0$ are 0 because $\lambda_{N}=\lambda_{N}^{*}=0$). The integrals in the second line can be evaluated, by integration by parts, as
@@
$\begin{split}\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r&=\left[-(T_{i\pm 1}^{*}+r)^{m}\frac{e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}}{\lambda_{i}^{*}+\mu_{i}^{*}}\right]_{0}^{\infty}+m\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m-1}\frac{e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathrm{d}r\\ &=\frac{(T_{i\pm 1}^{*})^{m}}{\lambda_{i}^{*}+\mu_{i}^{*}}+\frac{m}{\lambda_{i}^{*}+\mu_{i}^{*}}\int_{0}^{\infty}(T_{i\pm 1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r,\end{split}$ (B4)
so $\mathbb{E}\left[(T_{i}^{*})^{m}\right]$ turns into
$\begin{split}&\mathbb{E}\left[(T_{i}^{*})^{m}\right]=\frac{\lambda_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i+1}^{*})^{m}\right]+\frac{\mu_{i}^{*}}{\lambda_{i}^{*}+\mu_{i}^{*}}\mathbb{E}\left[(T_{i-1}^{*})^{m}\right]\\ &\phantom{=}+\frac{m}{\lambda_{i}^{*}+\mu_{i}^{*}}\left\{\lambda_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i+1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]+\mu_{i}^{*}\mathbb{E}\left[\int_{0}^{\infty}(T_{i-1}^{*}+r)^{m-1}e^{-(\lambda_{i}^{*}+\mu_{i}^{*})r}\mathrm{d}r\right]\right\}.\end{split}$ (B5)
@@
Since the expression in curly brackets equals $\mathbb{E}\left[(T_{i}^{*})^{(m-1)}\right]$ by (B3), we obtain the first-step equation
$(\lambda_{i}^{*}+\mu_{i}^{*})\mathbb{E}\left[(T_{i}^{*})^{m}\right]=\lambda_{i}^{*}\mathbb{E}\left[(T_{i+1}^{*})^{m}\right]+\mu_{i}^{*}\mathbb{E}\left[(T_{i-1}^{*})^{m}\right]+m\mathbb{E}\left[(T_{i}^{*})^{(m-1)}\right],\quad m\geq 1,i\in[N],$ (B6)
with boundary conditions $\mathbb{E}\left[(T_{0}^{*})^{m}\right]=0$ for all $m\geq 1$ and $\mathbb{E}\left[(T_{i}^{*})^{0}\right]=1$ for all $i\in[N]$.
@@
Since (B6) expresses $m$th moments in terms of $(m-1)$st moments, we can now generalise Norden (1982) and recursively solve (B6) from lower to higher moments. To this end, let $\boldsymbol{b}^{(m)}\mathrel{\mathop{:}}=\left(\mathbb{E}\left[(T_{1}^{*})^{m}\right],\cdots,\mathbb{E}\left[(T_{N}^{*})^{m}\right]\right)^{\top}$ for $m\geq 0$; note that $\boldsymbol{b}^{(0)}=\boldsymbol{1}$, a vector whose components are all one. Then (B6) reads
$\boldsymbol{A}^{*}\boldsymbol{b}^{(m)}=m\boldsymbol{b}^{(m-1)},\quad m>0,$ (B7)
where $\boldsymbol{A}^{*}=(A^{*}_{ij})_{i,j\in[N]}$ is an $N\times N$ matrix with elements
${A}^{*}_{ij}=\begin{cases}\lambda_{i}^{*}+\mu_{i}^{*},&j=i,\\ -\lambda_{i}^{*},&j=i+1,\\ -\mu_{i}^{*},&j=i-1,\\ 0,&\text{otherwise}.\end{cases}$ (B8)
@@
It is also clear from (B2) that $\boldsymbol{A}^{*}=\boldsymbol{F}^{-1}\boldsymbol{A}$, where $\boldsymbol{F}$ is the (invertible) $N\times N$ diagonal matrix $\boldsymbol{F}=\textrm{diag}\left[f(1),\cdots,f(N)\right]$ and $\boldsymbol{A}=(A_{ij})_{i,j\in[N]}$ is defined as $\boldsymbol{A}^{*}$ with $\lambda_{i}^{*}$ and $\mu_{i}^{*}$ replaced by $\lambda_{i}^{*}$ and $\mu_{i}^{*}$, respectively. Therefore, (B7) turns into
$\left(\boldsymbol{F}^{-1}\boldsymbol{A}\right)\boldsymbol{b}^{(m)}=m\boldsymbol{b}^{(m-1)}$ (B9)
or, equivalently,
@@
From the standard theory of Markov chains (Kemeny and Snell 1960, Chap. III; see also Norden 1982, p. 694), we know that $\boldsymbol{A}$ is invertible and $\boldsymbol{A}^{-1}$ has elements $\left(\boldsymbol{A}^{-1}\right)_{ij}=\zeta_{ij}$ with $\zeta_{ij}$ of (22); so (24) is an immediate consequence.
C A derivation of the first two moments of the stationary tree length, $L_{n}$

### v2

@@ Acknowledgment
We are greatly indebted to Anton Wakolbinger (Frankfurt) for enlightening discussions about Muller’s ratchet with tournament selection and many other conceptual and technical aspects of the paper, including a more elegant proof of Proposition A.1. We also thank Tom Britton (Stockholm) for a helpful conversation about the connection to epidemiology. Two anonymous reviewers dedicated a lot of time and expertise to the manuscript and made invaluable suggestions to improve it. J.Y.W. received funding from MEXT/JSPS KAKENHI Grant Numbers JP21K03357 and JP24H00001. H.O. acknowledges the support from the "Evolutionary Studies of Complex Adaptive Systems" Research Grant from the Research Center for Integrative Evolutionary Science, SOKENDAI. Y. K. received funding from MEXT/JSPS KAKENHI Grant Numbers 25K06712 and 25K00199. E.B. received funding from the Deutsche Forschungsgemeinschaft, Germany (DFG, German Research Foundation) — Project-ID 317210226 — SFB 1283.
@@ A Sojourn times of the birth-death process
Consider our general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $N]_{0}$ with unique absorbing state 0 and birth and death rates $\lambda_{j}$ and $\mu_{j}$ with $\lambda_{0}=0$, $\lambda_{j}\geq 0$ for $j\in[N-1]$, and $\mu_{j}>0$ for $j\in[N]$, complemented by $\mu_{0}=\lambda_{N}=0$.
@@
Let $Z(0)=i\in[N]$, let $T_{\ell}$ be the time where the process hits state $\ell\;(1\leq\ell\leq i)$ for the first time, and let $S_{j}^{(\ell)}$ be the total sojourn time in state $j\;(\ell\leq j\leq N)$ in the interval $T_{\ell},T_{\ell-1}]$. We then have
$\mathbb{E}\big[S_{j}^{(\ell)}\mid Z(0)=i\big]=\frac{\lambda_{\ell}\cdots\lambda_{j-1}}{\mu_{\ell}\cdots\mu_{j}}=\mathrel{\mathop{:}}\eta_{\ell j}$
independently of $i$, where the empty product is 1.
@@
The $T_{\ell}$ are finite almost surely due to the almost sure absorption of $(Z(t))_{t\geqslant 0}$ in 0. If $\lambda_{\ell}\cdots\lambda_{j-1}=0$, the claim is trivially true. We therefore assume $\lambda_{\ell}\cdots\lambda_{j-1}>0$. Let $V_{j}$ be the number of visits the chain pays to state $j$ in $T_{\ell},T_{\ell-1}]$ (this includes the initial visit if $j=\ell$). Each such visit has a mean duration of $1/(\lambda_{j}+\mu_{j})$. By Wald’s identity, we have
$\mathbb{E}\big[S_{j}^{(\ell)}\mid Z(0)=i\big]=\frac{\mathbb{E}\left[V_{j}\right]}{\lambda_{j}+\mu_{j}}.$ (A1)
To compute $\mathbb{E}\left[V_{j}\right]$, we define a discrete-time birth-death process $(Z^{\prime}(n))_{n\geq 0}$ on $\{\ell-1,\ell,\dots,N\}$ with $Z^{\prime}(0)=\ell-1$ and transition probabilities
$p_{\ell-1,\ell}=1,\quad\text{together with }\,p_{k,k-1}=\frac{\mu_{k}}{\lambda_{k}+\mu_{k}}\quad\text{and }\,p_{k,k+1}=\frac{\lambda_{k}}{\lambda_{k}+\mu_{k}}\quad\text{for }\,\ell\leq k\leq N.$ (A2)
Then $T^{\prime}_{\ell}=1$ is the first time where $(Z^{\prime}(n))_{n\geq 0}$ reaches $\ell$, and we denote by $T^{\prime}_{\ell-1}$ the time of its first visit to $\ell-1$. Clearly, $(Z^{\prime}(n))_{n\geq 0}$ restricted to $T^{\prime}_{\ell},T^{\prime}_{\ell-1}]$ has the same law as the embedded jump process of $(Z(t))_{t\geq 0}$ restricted to $T_{\ell},T_{\ell-1}]$, so $V^{\prime}_{j}$, the number of visits of $(Z^{\prime}(n))_{n\geq 0}$ to $j$ in $T^{\prime}_{\ell},T^{\prime}_{\ell-1}]$, has the same law as $V_{j}$, and, in particular, $\mathbb{E}\left[V_{j}\right]=\mathbb{E}\left[V^{\prime}_{j}\right]$. We now complement the $V_{j}^{\prime}$ ($\ell\leq j\leq N$) by $V^{\prime}_{\ell-1}\mathrel{\mathop{:}}=1$ (for the initial visit of $(Z^{\prime}(n))_{n\geq 0}$ to $\ell-1$) and note that, for $\ell\leq j\leq N$, the final step of $(Z^{\prime}(n))_{n\geq 0}$ to $\ell-1$ does not count into $V_{j}^{\prime}$. This gives $V^{\prime}_{k}=\sum_{n=0}^{T_{\ell-1}-1}\boldsymbol{1}_{\{Z(n)=k\}}$ for $\ell-1\leq k\leq N$, and $v\mathrel{\mathop{:}}=(\mathbb{E}\left[V^{\prime}_{\ell-1}\right],\ldots,\mathbb{E}\left[V^{\prime}_{N}\right])$ constitutes an invariant measure of $(Z^{\prime}(n))_{n\geq 0}$, see Norris (1997, Thm. 1.7.5). On the other hand, $(Z^{\prime}(n))_{n\geq 0}$ is recurrent and has a unique stationary distribution, $u=(u_{\ell-1},u_{\ell},\ldots,u_{N})$ say; it satisfies the detailed balance condition $u_{k}p_{k,k+1}=u_{k+1}p_{k+1,k}$ for $\ell-1\leq k<N$. By Thm. 1.7.6 of Norris (1997), $v$ is proportional to $u$, so
$\mathbb{E}\left[V^{\prime}_{j}\right]=\frac{\mathbb{E}\left[V^{\prime}_{j}\right]}{\mathbb{E}\left[V^{\prime}_{\ell-1}\right]}=\frac{u_{j}}{u_{\ell-1}}=\frac{p_{\ell-1,\ell}\cdots p_{j-1,j}}{p_{\ell,\ell-1}\cdots p_{j,j-1}}=\frac{\lambda_{\ell}\cdots\lambda_{j-1}}{\mu_{\ell}\cdots\mu_{j}}(\lambda_{j}+\mu_{j}),$ (A3)
where the first equality comes from $V^{\prime}_{\ell-1}=1$, the second-last from detailed balance, and the last one from (A2). This, combined with $\mathbb{E}(V_{j})=\mathbb{E}(V_{j}^{\prime})$ and (A1), completes the proof. ∎
@@ B A derivation of the moments of the path integrals of birth-death processes
The moments (20) of the path integral (16) are usually derived via Laplace transforms (see, for example, Goel and Richter-Dyn (1974, App. D)). Here we take an alternative route that bypasses Laplace transforms and derives a first-step equation in a direct way, thus validating (62) derived in a heuristic way by Norden (1982) via the backward equation. Remember that our birth-death process $(Z(t))_{t\geqslant 0}$ is restricted to $N]_{0}$ with birth and death rates $\lambda_{j}\geq 0$ and $\mu_{j}>0$ for $j\in[N]$, $\lambda_{0}=0$, and the convention $\mu_{0}=\lambda_{N}=0$.
@@
Let us now calculate the $m$-th moment of our path integral $\mathcal{I}_{i}(f)$ of (16), following the standard first-step approach outlined, for example, in Pinsky and Karlin (2011, Sec. 3.4). Suppose that we are currently in state $i\in[N]$. The sojourn time in this state, whose realisation is denoted by $r$ below, follows the exponential distribution with parameter $\lambda_{i}+\mu_{i}$. This time period contributes $rf(i)$ to the path integral. After the transition away from $i$, the further contribution to the path integral until absorption is distributed as $\mathcal{I}_{i+1}(f)$ or $\mathcal{I}_{i-1}(f)$, depending on whether it moves from $i$ to $i+1$ (probability $\lambda_{i}/(\lambda_{i}+\mu_{i})$) or i1i-1 (probability $\mu_{i}/(\lambda_{i}+\mu_{i})$). Thus we obtain
@@
$\begin{split}&\mathbb{E}\left[(\mathcal{I}_{i}(f))^{m}\right]\\ &=\int_{0}^{\infty}\left\{\frac{\lambda_{i}}{\lambda_{i}+\mu_{i}}\mathbb{E}\left[\{\mathcal{I}_{i+1}(f)+rf(i)\}^{m}\right]+\frac{\mu_{i}}{\lambda_{i}+\mu_{i}}\mathbb{E}\left[\{\mathcal{I}_{i-1}(f)+rf(i)\}^{m}\right]\right\}(\lambda_{i}+\mu_{i})e^{-(\lambda_{i}+\mu_{i})r}\mathrm{d}r\\ &=\lambda_{i}\mathbb{E}\left[\int_{0}^{\infty}\{\mathcal{I}_{i+1}(f)+rf(i)\}^{m}e^{-(\lambda_{i}+\mu_{i})r}\mathrm{d}r\right]+\mu_{i}\mathbb{E}\left[\int_{0}^{\infty}\{\mathcal{I}_{i-1}(f)+rf(i)\}^{m}e^{-(\lambda_{i}+\mu_{i})r}\mathrm{d}r\right]\end{split}$ (B1)
@@
(products containing the factor $\lambda_{N}$ are 0). The integrals in the second line can be evaluated by integration by parts as
@@
$\begin{split}&\int_{0}^{\infty}\{\mathcal{I}_{i\pm 1}(f)+rf(i)\}^{m}e^{-(\lambda_{i}+\mu_{i})r}\mathrm{d}r\\ &=\left[-\{\mathcal{I}_{i\pm 1}(f)+rf(i)\}^{m}\frac{e^{-(\lambda_{i}+\mu_{i})r}}{\lambda_{i}+\mu_{i}}\right]_{0}^{\infty}+mf(i)\int_{0}^{\infty}\{\mathcal{I}_{i\pm 1}(f)+rf(i)\}^{m-1}\frac{e^{-(\lambda_{i}+\mu_{i})r}}{\lambda_{i}+\mu_{i}}\mathrm{d}r\\ &=\frac{(\mathcal{I}_{i\pm 1}(f))^{m}}{\lambda_{i}+\mu_{i}}+\frac{mf(i)}{\lambda_{i}+\mu_{i}}\int_{0}^{\infty}\{\mathcal{I}_{i\pm 1}(f)+rf(i)\}^{m-1}e^{-(\lambda_{i}+\mu_{i})r}\mathrm{d}r,\end{split}$ (B2)
so $\mathbb{E}\left[(\mathcal{I}_{i}(f))^{m}\right]$ turns into
[…]
@@
Since the expression in curly brackets equals $\mathbb{E}\left[(\mathcal{I}_{i}(f))^{m-1}\right]$ by (B1), we obtain the first-step equation
$(\lambda_{i}+\mu_{i})\mathbb{E}\left[(\mathcal{I}_{i}(f))^{m}\right]=\lambda_{i}\mathbb{E}\left[(\mathcal{I}_{i+1}(f))^{m}\right]+\mu_{i}\mathbb{E}\left[(\mathcal{I}_{i-1}(f))^{m}\right]+mf(i)\mathbb{E}\left[(\mathcal{I}_{i}(f))^{m-1}\right],\;m\geq 1,\,i\in[N],$ (B4)
with boundary conditions $\mathbb{E}\left[(\mathcal{I}_{0}(f))^{m}\right]=0$ for all $m\geq 1$ and $\mathbb{E}\left[(\mathcal{I}_{i}(f))^{0}\right]=1$ for all $i\in[N]$.
@@
Since (B4) expresses $m$-th moments in terms of (m1)(m-1)-st moments, we can now generalise Norden (1982) and recursively solve (B4) from lower to higher moments. To this end, let $\boldsymbol{b}^{(m)}\mathrel{\mathop{:}}=\left(\mathbb{E}\left[(\mathcal{I}_{1}(f))^{m}\right],\cdots,\mathbb{E}\left[(\mathcal{I}_{N}(f))^{m}\right]\right)^{\top}$ for $m\geq 0$; note that $\boldsymbol{b}^{(0)}=\boldsymbol{1}$, a vector whose components are all one. Then (B4) reads
$\boldsymbol{A}\boldsymbol{b}^{(m)}=m\boldsymbol{F}\boldsymbol{b}^{(m-1)},\quad m>0,$ (B5)
where $\boldsymbol{A}=(A_{ij})_{i,j\in[N]}$ is an $N\times N$ matrix with elements
${A}_{ij}=\begin{cases}\lambda_{i}+\mu_{i},&j=i,\\ -\lambda_{i},&j=i+1,\\ -\mu_{i},&j=i-1,\\ 0,&\text{otherwise},\end{cases}$ (B6)
@@
and $\boldsymbol{F}$ is the $N\times N$ diagonal matrix $\boldsymbol{F}=\textrm{diag}\left[f(1),\cdots,f(N)\right]$. Therefore, from (B5) we immediately obtain
$\boldsymbol{b}^{(m)}=m\left(\boldsymbol{A}^{-1}\boldsymbol{F}\right)\boldsymbol{b}^{(m-1)},$ (B7)
which leads to
@@
From the standard theory of Markov chains (Kemeny and Snell 1960, Chap. III; see also Norden 1982, p. 694), we know that $\boldsymbol{A}$ is invertible and $\boldsymbol{A}^{-1}$ has elements $\left(\boldsymbol{A}^{-1}\right)_{ij}=\zeta_{ij}$ with $\zeta_{ij}$ of (18) and (19); so (20) is an immediate consequence.
C Derivation of the first two moments of $L_{n}$